# Unleashing the Power of Task-Specific Directions in Parameter Efficient Fine-tuning

### v3

@@ E.2 More Results for the Alignment on the Directions of LoRA-Dash
$\Delta\mathbf{W}_{all}\cap{\rm LTSD}$ 1.00 0.75 1.00 0.74 1.00 1.00 1.00 0.99
${\rm TSD}\cap{\rm LTSD}$ 1.00 1.00 1.00 0.78 0.75 0.75 0.50 0.73
${\rm TSD}\cap\Delta\mathbf{W}_{all}$ 1.00 0.75 1.00 0.83 0.75 0.75 0.50 0.76
Table 22: Alignment on the directions of LoRA-Dash with TSDs. We report the percentage from the self-attention value (v) projection of the first, middle and last layer of two models, as well as the average percentage of all q, k, and v modules.
@@
$\Delta\mathbf{W}_{all}\cap{\rm LTSD}$ 0.75 0.75 0.75 0.75 0.75 1.00 1.00 0.97
${\rm TSD}\cap{\rm LTSD}$ 0.75 0.75 0.75 0.77 0.75 0.75 1.00 0.78
${\rm TSD}\cap\Delta\mathbf{W}_{all}$ 1.00 1.00 1.00 0.81 0.75 1.00 1.00 0.82
E.3 More Results for Ablation Study
We here provide more results when fine-tuning DeBERTaV3-large on the ablation study in Fig. 10(b)-(c) in Sec. 6.4. The experimental results observed align well with the trends analyzed in Sec. 6.4, indicating consistency between our theoretical insights and practical outcomes.
Figure 10: More results on CoLA when fine-tuning DeBERTaV3-large for ablation study of LoRA-Dash. a. Ablation study on the directions for dash. Beyond the top, random, and bottom directions, we also explored adjusting all directions and a combined adjustment of both bottom and top directions. b. Ablation study on the length of pre-launch phase $t$. c. Ablation study on the number of directions $s$ for dash.
@@ E.3 More Results for Ablation Study
We here further discuss the impact of selecting different dash directions. The effectiveness of choosing top, bottom, or random directions has been a topic frequently mentioned in various studies. However, we find differing conclusions in different works, which complicates the understanding. For instance, Meng et al. (2024) and Zhang & Pilanci (2024) suggest that selecting top directions yields the best results, while Hameed et al. (2024) advocates for random selections, and reference Wang et al. (2024) supports choosing the bottom. This inconsistency is puzzling and is a matter we aim to explore further through our research.
Indeed, fine-tuning leverages a model’s broad pre-trained knowledge base, adapting it to learn specific, nuanced details relevant to particular downstream tasks. The essence of TSDs lies in their representation of the singular directions where the original model weights exhibit the most significant changes during adaptation to new tasks. Intuitively, these TSDs, being the most altered, should indeed offer the most impactful directions for adjustment during fine-tuning, potentially yielding the best performance.
Analyzing the properties of TSDs reveals that they typically reside along the directions corresponding to smaller, yet not the smallest, singular values. If we choose to dash in the directions associated with the smallest singular values—the “bottom” directions—it’s plausible that some of these might overlap with the TSDs. Hence, dashing in the bottom directions might produce results inferior to directly using the TSDs.
As for the top singular directions, which encapsulate the most significant information from pre-training, these generally involve more generalized knowledge. Since such information is less task-specific, adjusting these directions could potentially degrade performance rather than enhance it, especially if the task at hand requires nuanced understanding that diverges from the general pre-training data.
Lastly, a random selection of directions introduces unpredictability in outcomes. However, there is a chance that such random choices might inadvertently select appropriate directions or even TSDs, thus performing better than top selections but generally worse than bottom choices due to the randomness. This method could strike a balance between leveraging underlying model strengths and adapting to new specifics, albeit inconsistently.
@@
In addition to the standard experimental setup, we conducted two additional experiments: adjusting the coordinate value changes of all directions or both top and bottom directions collectively. Neither of these configurations resulted in improved outcomes. This aligns with our analysis in Sec. 6.4, which suggested that the inclusion of irrelevant directions could degrade training performance due to increased noise in the model updates.
E.4 Are TSDs “Task-specific”?
We sought to verify whether TSDs are genuinely task-specific, that is, whether the most critical directions differ for each task. To investigate this, we employed a prototypical example: the subject-driven generation task. In this setting, we tracked the TSDs captured by LoRA-Dash when different subjects served as targets and analyzed the relationships between these TSDs. The results are presented in the Fig. 11 (Please refer to the caption of Fig. 11 for detailed information).
Figure 11: The results of the TSDs similarity between different tasks (subjects). For the first row: Images from different subjects. The second and third rows contain results extracted from two different layers of U-net Ronneberger et al. (2015). Each subplot illustrates the similarity of TSDs captured for two different subjects. The title of each subplot indicates the respective subjects, with the directional order of the first subject represented by columns and the second by rows. The deep blue blocks highlight where the directions for the two subjects coincide.
E.5 Convergence Curve
We present loss curves for both LoRA-Dash and LoRA when fine-tuning LLaMA-7B. We ensure consistency in all training hyper-parameters, and the results are shown in Fig. 12(a). We can observe that during the pre-launch phase, the loss trajectories of LoRA-Dash and LoRA are aligned. However, as the training progresses into the dash phase, LoRA-Dash consistently exhibits a slightly lower loss compared to standard LoRA.
One may concern that a lower training loss could lead to overfitting. Therefore, we also tracked the validation loss throughout the training process, with the results presented in Fig. 12(b). It is clear that the validation loss for LoRA-Dash is much lower than that for LoRA. This suggests that LoRA-Dash not only avoids overfitting but also enhances its learning efficacy by effectively incorporating task-specific knowledge.
Table 23: Results on commonsense reasoning tasks of LoRA-Dash compared with other methods. For LoRA derivatives, we report the parameter gains over LoRA.
Method Params BoolQ PIQA SIQA HellaS. WinoG. ARC-e ARC-c OBQA Avg.
ChatGPT - 73.1 85.4 68.5 78.5 66.1 89.8 79.9 74.8 77.0
Fine-tuning LLaMA-7B
Fully FT 100% 69.9 84.2 78.9 92.3 83.3 86.6 72.8 83.4 81.4
Prefix 0.11% 64.3 76.8 73.9 42.1 72.1 72.9 54.0 60.6 64.6
Series 0.99% 63.0 79.2 76.3 67.9 75.7 74.5 57.1 72.4 70.8
Parallel 3.54% 67.9 76.4 78.8 69.8 78.9 73.7 57.3 75.2 72.2
LoRAr=4 0.10% 2.3 46.1 18.3 19.7 55.2 65.4 51.9 57.0 39.5
AdaLoRAr=4 + 0.6k 66.1 78.1 74.3 34.0 74.4 76.7 57.5 71.2 66.5
FLoRAr=4 + 2.6k 67.2 78.0 72.9 65.4 73.8 73.8 55.3 71.8 69.8
DoRAr=4 + 877k 51.3 42.2 77.8 25.4 78.8 78.7 62.5 78.6 61.9
LoRA-Dashr=4 + 1.3k 65.2 79.9 78.3 82.8 77.1 78.6 65.4 78.4 75.7
LoRAr=32 0.83% 68.9 80.7 77.4 78.1 78.8 77.8 61.3 74.8 74.7
AdaLoRAr=32 + 5.1k 69.1 82.2 77.2 78.3 78.2 79.7 61.9 77.2 75.5
FLoRAr=32 + 164k 66.4 81.3 77.1 75.6 77.1 77.2 62.4 77.6 74.3
DoRAr=32 + 877k 69.7 83.4 78.6 87.2 81.0 81.9 66.2 79.2 78.4
LoRA-Dashr=32 + 1.3k 69.9 82.8 78.6 84.9 81.6 82.3 66.5 80.8 78.4
Fine-tuning LLaMA3-8B
LoRAr=16 0.35% 72.3 86.7 79.3 93.5 84.8 87.7 75.7 82.8 82.8
AdaLoRAr=16 + 2.6k 90.4 85.0 76.7 79.1 83.3 86.4 75.1 75.4 81.4
FLoRAr=16 + 41k 90.2 84.2 79.9 79.3 85.1 86.7 74.8 93.9 84.2
LoRA-Dashr=16 + 1.3k 74.8 88.0 80.6 95.2 85.6 89.0 77.4 84.8 84.4
LoRAr=32 0.70% 70.8 85.2 79.9 91.7 84.3 84.2 71.2 79.0 80.8
PISSAr=32 + 0 67.1 81.1 77.2 83.6 78.9 77.7 63.2 74.6 75.4
MiLoRAr=32 + 0 68.8 86.7 77.2 92.9 85.6 86.8 75.5 81.8 81.9
DoRAr=32 + 784k 74.6 89.3 79.9 95.5 85.6 90.5 80.4 85.8 85.2
LoRA-Dashr=32 + 1.3k 75.3 88.5 80.2 95.7 86.8 90.7 80.2 85.6 85.4
Table 24: Results with DeBERTaV3-base fine-tuned on GLUE development set. “FT” represents fully fine-tuning.
Method % Params MNLI SST-2 CoLA QQP QNLI RTE MRPC STS-B All
Acc Acc Mcc Acc Acc Acc Acc Corr Avg.
FT 100% 89.90 95.63 69.19 91.87 94.03 83.75 90.20 91.60 88.27
(IA)3 0.03% 89.44 95.52 67.01 89.01 91.80 79.42 88.23 90.79 86.40
SSL 0.02% 88.35 95.07 66.64 88.19 90.10 82.31 88.68 90.13 86.18
SSB 0.05% 89.86 95.53 67.82 89.87 93.41 83.75 88.72 90.94 87.49
BitFit 0.05% 89.37 94.84 66.96 88.41 92.24 78.80 87.75 91.35 86.21
Series 0.17% 90.10 95.41 67.65 91.19 93.52 83.39 89.25 91.31 87.73
PAdapter 0.16% 89.89 94.72 69.06 91.05 93.87 84.48 89.71 91.38 88.02
LoRA 0.18% 90.03 93.92 69.15 90.61 93.37 87.01 90.19 90.75 88.13
AdaLoRA 0.18% 90.66 95.80 70.04 91.78 94.49 87.36 90.44 91.63 88.86
FLoRA 0.18% 90.60 96.00 70.20 91.40 94.46 88.09 90.93 91.96 89.21
DoRA 0.22% 90.21 94.38 69.33 90.84 93.26 86.94 90.19 91.34 88.31
LoRA-Dash 0.18% 90.14 95.42 72.41 91.65 94.36 89.89 91.67 91.64 89.65
E.6 Comparison with Other Methods
In the main text, we only mainly compared the results of LoRA-Dash and LoRA because LoRA-Dash is designed to further harness the potential of TSDs captured by LoRA. Here, we further compare the performance of LoRA-Dash with other methods.
Specifically, we compare BitFit Zaken et al. (2021), Prompt Learning (Prefix) Li & Liang (2021), Series Adapter (Series) Houlsby et al. (2019), Adapter proposed by Pfeiffer et al. (2020) (PAdapter), Parallel Adapter (Parallel) He et al. (2021a), (IA)3 Liu et al. (2022), SSL Si et al. (2024b), SSB Si et al. (2024b), AdaLoRA Zhang et al. (2022), PISSA Meng et al. (2024), MiLoRA Wang et al. (2024), FLoRA Si et al. (2024a) and DoRA Liu et al. (2024). We follow the settings detailed in our paper, and the results are shown in Tables. 23-24.
The results clearly show that even compared to state-of-the-art (SOTA) methods like DoRA and FLoRA, LoRA-Dash holds its ground impressively. Additionally, two phenomena are noteworthy:
  • When fine-tuning LLaMA-7B, at a rank of 4, all methods except LoRA-Dash perform poorly under low parameter budgets, while LoRA-Dash achieves results comparable to those with higher parameter settings. This further highlights the superiority of LoRA-Dash, which stimulates performance on downstream tasks by unleashing the potential of TSDs.
  • Compared with LoRA derivatives, LoRA-Dash introduces significantly fewer additional parameters, such as nearly 130 times fewer than FLoRA and 700 times fewer than DoRA at rank 8 when fine-tuning LLaMA-7B, yet achieves better or comparable results.
Revisiting the objective of Parameter-Efficient Fine-Tuning, the key question is whether it is possible to achieve performance comparable to fully fine-tuning while training with significantly fewer parameters. Through extensive experimentation and comparison with other methods, LoRA-Dash maintains excellent performance even with a minimal parameter budget, where most methods would be constrained by the amount of parameters available. This demonstrates that LoRA-Dash significantly surpasses other methods in terms of parameter utilization efficiency. Therefore, we conclude that LoRA-Dash exemplifies an ideal realization of PEFT goals.
Reflecting on the entirety of this research, we arrive at a final conclusion:
“LoRA-Dash epitomizes the essence of parameter efficiency, maximizing the efficacy of each parameter deployed. With significantly fewer parameters trained, LoRA-Dash sustains or even exceeds the performance set by fully fine-tuning.”
This marks the end of our current exploration, but it is merely a pause in the ongoing quest for deeper understanding. The path ahead remains rich with potential, inviting further inquiry into the intricate dance of knowledge…
The End

### v4

@@ 5.5 Convergence Curve
DTSDs $\cap$ LTSDs 1.00 1.00 1.00 0.76 1.00 1.00 1.00 0.99
TSDs $\cap$ LTSDs 1.00 1.00 1.00 0.78 0.75 0.75 0.50 0.73
TSDs $\cap$ DTSDs 1.00 1.00 1.00 0.85 0.75 0.75 0.75 0.82
Model LLaMA-7B ($r=16$) DeBERTaV3-large ($r=8$)
DTSDs $\cap$ LTSDs 1.00 1.00 0.75 0.75 1.00 1.00 1.00 0.98
TSDs $\cap$ LTSDs 0.75 0.75 1.00 0.74 1.00 1.00 1.00 0.84
TSDs $\cap$ DTSDs 1.00 0.75 1.00 0.86 1.00 1.00 1.00 0.91
@@
DTSDs $\cap$ LTSDs 1.00 0.75 0.75 0.76 0.75 1.00 1.00 0.96
TSDs $\cap$ LTSDs 0.75 0.75 0.75 0.77 0.75 0.75 1.00 0.82
TSDs $\cap$ DTSDs 1.00 1.00 1.00 0.81 0.75 1.00 1.00 0.79
6 Analysis of Task-specific Directions
In this section, we conduct extensive experiments to better understand task-specific directions, with the method LoRA-TSD for illustration. We first explore the relationships between the LTSDs, the final directions after the dash phase, and TSDs to determine whether LoRA-TSD can amplify the information of TSDs (Sec. 6.1) and, if so, to what extent (Sec. 6.2). We then explore whether TSD are “task-specific” (Sec. 6.3), and verify whether the information in TSDs can help enhance the effectiveness of other methods (Sec. 6.4). Finally, We provide more detailed discussions on the task-specific directions.
6.1 How Do TSDs Align with the Final Directions?
We examine the alignment between the LTSDs identified during the pre-launch phase, the directions with highest change rates pinpointed by the final weight change of LoRA-TSD (denoted as “DTSDs”, Delta TSDs), and the true TSDs derived from $\mathbf{W}^{*}$. Following the settings in Secs. 3.4, we report the percentage that:
  • How many of the top 4 DTSDs directions are contained in $s$ LTSDs [^4] Here since $s=8$, the setting is equal to that in Sec. 3.4.. We denote this percentage as “DTSDs $\cap$ LTSDs ” for further convenience.
The first is to explore the consistency between the directions selected during the pre-launch phase and those showing the greatest change after the final training, and the second and third aim to validate the consistency between the directions chosen during the pre-launch phase and the final trained with the true TSDs. We conducted tests on LLaMA-7B [4] and DeBERTaV3-large [3], each with three ranks of LoRA-TSD, and the results are shown in Table. V. It is obvious that LoRA-TSD indeed captures a great proportion of TSDs’ information. Moreover, we can also obtain two significant findings: First, the LTSDs and DTSDs of LoRA-TSD align closely, which indicates that the directions amplified by LoRA-TSD finally are essentially consistent with those selected during the pre-launch phase. Therefore, since LTSDs contain substantial information about TSDs, it is not surprising that the final DTSDs consistently include information about the true TSDs. Second, when compared with the recall results in Fig. 3 where the end of each line indicates the final percentage of TSDs $\cap$ DTSDs of LoRA, it becomes evident that LoRA-TSD’s trained weights capture a greater proportion of TSD information than vanilla LoRA.
6.2 How much can TSDs be Amplified?
Since LoRA-TSD captures the information of TSDs, we further explore to what extent LoRA-TSD can amplify the information of these directions. To quantify this amplification, we define an amplification factor that measures how significantly the task-specific features are enhanced. We first project the pretrained weights $\mathbf{W}$ and the weights merged after training, i.e., $\mathbf{W}_{res}+\Delta\mathbf{W}_{all}$, onto the LTSDs. Let $\bar{\mathbf{U}}=\begin{bmatrix}\bar{\mathbf{u}}_{1}&\ldots&\bar{\mathbf{u}}_{s}\end{bmatrix}$ and $\bar{\mathbf{V}}=\begin{bmatrix}\bar{\mathbf{v}}_{1}&\ldots&\bar{\mathbf{v}}_{s}\end{bmatrix}$, we can then define the amplification factor of $\Delta\mathbf{W}_{all}$ as the Frobenius norms of these two weights on these directions as $\frac{\|\bar{\mathbf{U}}^{\mathsf{T}}\mathbf{W}_{tsd}\bar{\mathbf{V}}\|_{F}}{\|\bar{\mathbf{U}}^{\mathsf{T}}\mathbf{W}\bar{\mathbf{V}}\|_{F}}$. Specifically, we also evaluate the separate impacts of $\Delta\mathbf{W}_{AB}$ and $\Delta\mathbf{W}_{dash}$ in the amplification, with the corresponding amplification factor being $\frac{\|\bar{\mathbf{U}}^{\mathsf{T}}(\mathbf{W}_{res}+\Delta\mathbf{W}_{AB})\bar{\mathbf{V}}\|_{F}}{\|\bar{\mathbf{U}}^{\mathsf{T}}\mathbf{W}\bar{\mathbf{V}}\|_{F}}$ and $\frac{\|\bar{\mathbf{U}}^{\mathsf{T}}(\mathbf{W}_{res}+\Delta\mathbf{W}_{dash})\bar{\mathbf{V}}\|_{F}}{\|\bar{\mathbf{U}}^{\mathsf{T}}\mathbf{W}\bar{\mathbf{V}}\|_{F}}$.
We conduct this experiment on LLaMA-7B with three ranks. The results shown in Table. VI demonstrate that LoRA-TSD significantly enhances the features associated with LTSDs under all test conditions. Notably, the feature amplification contributed by $\Delta\mathbf{W}_{dash}$ accounts for the majority of the overall enhancement, whereas the amplification from $\Delta\mathbf{W}_{AB}$ is relatively minor in comparison. LoRA-TSD effectively learns and amplifies the majority of TSD information with only $s$ parameters (i.e., $\Delta\sigma_{i}$ in $\Delta\mathbf{W}_{dash}$), reducing the learning burden on $\Delta\mathbf{W}_{AB}$ during fine-tuning. This efficiency also allows $\Delta\mathbf{W}_{AB}$ to potentially focus on learning other significant features, further optimizing the model’s performance on downstream tasks.
TABLE VI: Amplification on task-specific features of LoRA-TSD. We report the amplification factor from the self-attention key (k) projection of the first, middle, and last layers of LLaMA-7B, as well as the average of all q, k, and v modules.
Layer $r=8$ $r=16$ $r=32$
1st 285.12 45.32 280.05 635.22 173.58 595.75 45.50 15.62 40.20
16th 12.50 5.10 10.02 14.25 8.12 12.01 2.35 1.72 2.10
32nd 70.45 8.00 68.02 32.10 12.75 30.40 8.85 1.80 8.75
Avg. 23.56 7.88 22.37 33.52 13.82 30.04 5.57 2.71 5.35
Figure 9: The results of the TSDs similarity between different tasks (subjects). For the first row: Images from different subjects. The second and third rows contain results extracted from two different layers of U-net [49]. Each subplot illustrates the similarity of TSDs captured for two different subjects. The title of each subplot indicates the respective subjects, with the directional order of the first subject represented by columns and the second by rows. The deep blue blocks highlight where the directions for the two subjects coincide.
6.3 Are TSDs “Task-specific”?
We sought to verify whether TSDs are genuinely task-specific, that is, whether the most critical directions differ for each task. To investigate this, we employed a prototypical example: the subject-driven generation task. In this setting, we tracked the TSDs captured by LoRA-TSD when different subjects served as targets and analyzed the relationships between these TSDs. The results are presented in Fig. 9.
6.4 Can TSDs Enhance Other Methods’ Performance?
To investigate whether TSDs are beneficial for other PEFT methods, we extended our experiments to include AdaLoRA [12] and FLoRA [14], assessing the impact of leveraging TSDs within these frameworks. The results, as shown in Table. VII, clearly indicate that incorporating TSDs significantly enhances the performance of these methods as well. This underscores the pivotal role of TSDs in optimizing model behavior for downstream tasks, highlighting their universal applicability across various PEFT strategies.
TABLE VII: Results of commonsense reasoning tasks when fine-tuning LLaMA-7B.
ChatGPT - 73.1 85.4 68.5 78.5 66.1 89.8 79.9 74.8 77.0
AdaLoRAr=32 0.83 69.1 82.2 77.2 78.3 78.2 79.7 61.9 77.2 75.5
AdaLoRA-Dash 0.83 69.2 82.3 77.5 78.4 77.1 80.1 63.4 79.6 76.0
FLoRAr=32 0.83 66.4 81.3 77.1 75.6 77.1 77.2 62.4 77.6 74.3
FLoRA-Dash 0.83 69.8 81.9 78.0 83.3 79.6 79.1 62.7 79.4 76.7
6.5 Further Discussion on Directions for Dash
@@
We here further discuss the impact of selecting different dash directions. The effectiveness of choosing top, bottom, or random directions has been a topic frequently mentioned in various studies. However, we find differing conclusions in different works, which complicates the understanding. For instance, [22] and [50] suggest that selecting top directions yields the best results, while [23] advocates for random selections, and reference [24] supports choosing the bottom. This inconsistency warrants further exploration.
Fine-tuning adapts a model’s pre-trained knowledge to task-specific details. TSDs correspond to the directions with the most significant weight changes during adaptation. Intuitively, these should be the most impactful for fine-tuning. TSDs typically align with directions of moderate singular values, suggesting that dashing in bottom directions might overlap with TSDs but yield worse results. Top directions, containing more generalized pre-training knowledge, could degrade performance when fine-tuning for task-specific details. Random direction selection introduces unpredictability, occasionally matching TSDs but generally performing worse than bottom selections.
@@
In addition to the standard experimental setup, we conducted two additional experiments: adjusting the coordinate value changes of all directions or both top and bottom directions collectively in Fig. 7. Neither of these configurations resulted in improved outcomes. This aligns with our analysis in Sec. 5.3, which suggested that the inclusion of irrelevant directions could degrade training performance due to increased noise in the model updates.
7 Conclusion
In this paper, we revisit LoRA’s exploration of task-specific directions (TSDs), highlighting its inadequacies in understanding TSDs. We then provide a precise definition of TSD and delve into their properties to better understand their role in fine-tuning large language models. Building on this foundational knowledge, we introduce LoRA-Dash, LoRA-Init, and LoRA-TSD, designed to fully unleash the potential of TSDs. Through comprehensive experiments and in-depth analysis, we not only demonstrate the significant advantages of our methods over conventional methods, but also highlight the crucial importance of TSDs in achieving superior task-specific performance. By pushing the boundaries of parameter-efficient fine-tuning, we aim to inspire continued research and development in this vibrant area, transforming practices across diverse applications in natural language processing and beyond.